Projection Close 4/7
To do this evening/ was completed: Quiet night on our end, all in all.
Received the film schedule a couple minutes ago, so I will lock in with that tomorrow morning (I'm on admin tomorrow, Alec is at WB!)
Presentation Issues:
Preshow Clips/Playlist: Double checked the playlist for Return of the Living Dead, which played to a happy crowd tonight.
Got rid of old playlists and old content from over the weekend.
Trailer Updates:
Special Events/ AV set up: Took a closed captioning device out for T5. Back charging.
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): Ingested Miracle Mile to the LMS.
Marquee updated?:
Posters updated?:
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Trailer Updates:
Special Events/ AV set up: Took a closed captioning device out for T5. Back charging.
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): Ingested Miracle Mile to the LMS.
Ingested Sunset Boulevard to the LMS.
Outgoing/Shipping (content):
Print/video Inspection: SOF print came in, brought it down to T1. Inspected the print. In good condition despite some edge damage (particularly in reel 4). Color is great. Report is available, I have a few pics, but since it's SOF, I'll keep em until requested.
Asset tracking doc updated/checked?:
Maintenance/cleaning: Swept and mopped.
Cue sheets updated/checked (every shift): Yes
Outgoing/Shipping (content):
Print/video Inspection: SOF print came in, brought it down to T1. Inspected the print. In good condition despite some edge damage (particularly in reel 4). Color is great. Report is available, I have a few pics, but since it's SOF, I'll keep em until requested.
Asset tracking doc updated/checked?:
Maintenance/cleaning: Swept and mopped.
Cue sheets updated/checked (every shift): Yes
Marquee updated?:
Posters updated?:
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Comments
Post a Comment